OpenStudy (anonymous):

Solve the system by substitution -x-y-z=-8 -4x+4y+5z=7 2x+2z=4

OpenStudy (anonymous):

since we have an equation that only has to variables, we will need to get another equation to those two variables

OpenStudy (anonymous):

so we need to replace y, and the easiest way to find y= would be in the first equation

OpenStudy (anonymous):

so what does y=?

OpenStudy (anonymous):

I don't know...

OpenStudy (***[isuru]***):

from the first one ... -x-y-z=-8 -( x + y + z) = -8 x + y + z = 8 ---- (1) from the last one 2x+2z=4 2 ( x + z ) = 4 x + z = 2 ---- (2) now substitute x + z = 2 in ( 1) y + 2 = 8 y = 6 from (2) x + z = 2 z = 2 -x now take .. -4x+4y+5z=7 -4( x - y) + 5z = 7 substitute z = 2 -x here.. -4( x - y) + 5z = 7 -4( x - y) + 5(2 -x)= 7 -4x+4y + 10 - 5x = 7 -9x + 4y = 7 -10 = -3 substitute y = 6 -9x + 24 = -3 9x = 27 x = 3 from (2) z = 2 -x = 2 - 3 = -1 hope this will help ya!!

OpenStudy (anonymous):

thats okay -x-y-z=-8 so we want y alone so add x and y to both sides -y=-3+x+z then multiply by -1 go make y positive y=3-x-z

OpenStudy (anonymous):

so this is what we will substitute for y in the 2nd equation

OpenStudy (anonymous):

woah sorry not 3 8

OpenStudy (anonymous):

-4x+4y+5z=7 so lets input y -4x+4(8-x-z)+5z=7 distribute -4x+32-4x-4z+5z=7 then simplify -8x+z+32=7 subtract 32 from both sides -8x+z=-25

OpenStudy (anonymous):

so sorry

OpenStudy (anonymous):

anyway -8x+z=-25 2x+2z=4 divide by -2 to remove z -8x+z=-25 -x-z=-2 add the two equations -9x=-27 x=3

OpenStudy (anonymous):

lets reinput x into 2x+2z=4 2(3)+2z=4 simplify 2z+6=4 subtract -6 from both sides 2z=-2 divide by 2 z=-1

OpenStudy (anonymous):

so now we input x and z into the first equation -x-y-z=-8 -3-y-(-1)=-8 simplify -3-y+1=-8 -y-2=-8 -y=-6 y=6

OpenStudy (anonymous):

so x=3 y=6 z=-1

OpenStudy (anonymous):

Thank you!! Can you help me with another one?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Solve the system by elimination -2x+2y+3z=0 -2x-y+z=-3 2x+2z=4

OpenStudy (anonymous):

@insanitycat

OpenStudy (anonymous):

so lets get y from the second equation -2x-y+z=-3 so add 2x to both sides and subtract z -y=-3+2x-z now multiply by -1 to get y positive y=3-2x+z

OpenStudy (anonymous):

not lets substitute y into the first equation -2x+2y+3z=0 -2x+2(3-2x+z)+3z=0 distribute -2x+6-4x+2z+3z simplifly -6x+5z+6=0 subtract 6 from both sides -6x+5z=-6

OpenStudy (anonymous):

so now we have 2x+2z=4 lets multiply this equation by 3 -6x+5z=-6 6x+6z=12 -6x+5z=-6 add the two 11z=6 z=6/11

OpenStudy (anonymous):

lets input z into 2x+2z=4 2x+2(6/11)=4 2x+12/11=4 or 44/11 subtract 12/11 from both sides 2x=32/11 multiply by 1/2 x=16/11

OpenStudy (anonymous):

now lets input both into the 2nd equation -2x-y+z=-3 -2(16/11)-y+6/11=-3 simplify -32/11-y+6/11=-3 -y-26/11=-3 or -33/11 add 26/11 to both sides -y=-7/11 multiply by -1 y=7/11

OpenStudy (anonymous):

so x=16/11 y=7/11 and z=6/11

OpenStudy (anonymous):

Okay so what I have is 1. Get y from second equation -2x-y+z=-3 2. Add 2x to both sides and subtract z -y=3-2x+z 3. Multiply by -1 to get y positive y=3-2x+z 4. Substitute y into first equation -2x+2y+3z=0 becomes -2x+2(3-2x+z)+3z=0 5. Distribute -2x+6-4x+2z+3z 6. Simplify -6x+5z+6=0 7. Subtract 6 from both sides -6x+5z=-6

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

8. Multiply the equation 2x+2z=4 by 3 -6x+5z=6

OpenStudy (anonymous):

I actually messed up the equation. It's supposed to be Solve the system by elimination -2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5

OpenStudy (anonymous):

so lets elimate to get equations with two variables 2x+3y+3z=5 -2x+2y+3z=0 5y+6z=5 and 2x+3y+3z=5 -2x-y+z=-3 2y+5z=-2

OpenStudy (anonymous):

sorry 2y+4z=2 5y+6z=5 multiply this by 2 2y+4z=2 multiply this by -3 10y+12z=10 -6y-12z=-6 4y=4 y=1

OpenStudy (anonymous):

so substitute y into one of the equations ill use 2y+4z=2 2(1)+4z=2 4z+2=2 4z=0 z=0

OpenStudy (anonymous):

so lets subsitute y and z now 2x+3y+3z=5 2x+3(1)+3(0)=5 2x+3=5 2x=2 x=1

OpenStudy (anonymous):

so x=1 y=1 and z=0